# MSM_pseudo_CS

### v1

@@ 7 Breast cancer example (EORTC-trial 10854)
Table 5 summarizes the distribution of the state occupation at the inspection times. As expected, most patients were inspected while occupying state 0 (alive after surgery). We are interested in estimating the probability of a patient ever reaching state 5 given a previous occupation of state 1. Inference around this quantity is crucial in breast cancer studies (Vicini et al., 2003; Murata et al., 2023) since identifying patients at higher risk of distinct metastasis following local-regional relapse informs better treatment decisions to prevent or manage metastatic disease. This conditional quantity, denoted by $\Psi_{5|1}$, was estimated using both the FRE and PLE approaches.
The estimates and corresponding 95% bootstrap confidence intervals based on the FRE and PLE approach are $0.400\ (0.241,0.570)$ and $0.433\ (0.222,0.658)$, respectively. The estimation results from both proposed methods are comparable. To provide more context to the results, we conducted a purely marginal analysis (without conditioning on a previous visit to state 1) where we estimated the probability of ever reaching state 5. The resulting estimate is $0.050$. As expected, conditioning on the previous occupation of the relapse state significantly increases the probability of ever reaching state 5.
We further estimated $\Psi_{5|1}$ using the PLE approach on the original right-censored data. This yielded an estimate of $0.344$. Notably, the estimations based on the heavily censored cross-sectional data approximate the result obtained from the right-censored data, suggesting their potential as an alternative analysis approach.
Table 5: Distribution of state occupation at inspection based on the current status data from the EORTC-trial 10854.
State 0 1 2 3 4 5 6 7 8
Frequency 1941 119 122 66 93 57 29 30 336
Percent 0.695 0.043 0.044 0.024 0.033 0.020 0.010 0.011 0.120
@@ 8 Discussion
This paper carefully employs techniques from the framework of competing risks to construct nonparametric estimators of future state entry probabilities and future state entry distributions, predicated on the prior occupation of a state, utilizing current status multistate data. The cross-sectional nature of the censored event history data brings about additional complexities when attempting to construct nonparametric estimates of temporal functions within a multistate model. We tackle this problem by harnessing ideas from nonparametric regression modeling to formulate estimators of the counting process and at-risk sets that serve as foundational elements in constructing estimators of the probability functions. To estimate the conditional quantities in the context of current status data, we employed a method (FRE) that was only available for right-censored data, and we also introduced a novel method (based on PLE). Using data from a real-world application and simulation studies for illustrative purposes, we show that both estimation methods yield valid and comparable results.
In practice, each estimation method has its benefits and challenges. Constructing the estimator based on the FRE approach could appear challenging since it requires the calculation of the fractional weights. Although the PLE method may be easier to implement once the artificial states have been specified correctly, the FRE method may be preferred for calculations involving distant states since it produces more reliable estimates for such states.
The intuitive approaches we propose provide a more direct route for analyzing the severely limited event data since they utilize the averaging principle rather than maximizing a likelihood function. Suppose we consider a less complicated setting (survival model with two states); the nonparametric maximum likelihood estimator (NPMLE) of the survival function has undesirable properties since it is undefined over a set of regions. For the general multistate model with current status data, an NPMLE approach would be more problematic.
@@
Further, researchers analyzing current status multistate data may also be interested in assessing the impact of measured risk factors on the probability of occupying a future state given a prior visit to a particular state. Regression modeling based on jackknife pseudo-values (Andersen et al., 2003) presents a flexible approach for estimating covariate effects in event history analysis. Recently, Anyaso-Samuel et al. (2023) extended the pseudo-value approach to model the effect of covariates on state occupation probabilities with multistate current status data. By jackknifing any of the estimators proposed here and utilizing the guidelines provided in Anyaso-Samuel et al. (2023), researchers can easily conduct covariate inference for the future state entry probabilities.

### v2

@@ 8 Application to emulated current status data from a real-world study
We focused on estimating the probability of occupying state 5 (distant metastasis), conditional on a prior visit to state 1 (loco-regional recurrence only). This probability, denoted $\Psi_{5|1}$, is of clinical interest for identifying patients at increased risk of progression following recurrence [30, 22].
The estimates and corresponding 95% bootstrap confidence intervals based on the FRE and PLE approach are 0.400(0.241,0.570)0.400\ (0.241,0.570) and 0.433(0.222,0.658)0.433\ (0.222,0.658), respectively. The estimation results from both proposed methods are comparable. For context, a marginal (unconditional) analysis estimated the probability of occupying state 5 as only 0.050, emphasizing how conditioning on prior recurrence substantially changes the estimated risk. We also applied the PLE approach to the original right-censored data, yielding an estimate of 0.344 with a 95% CI of (0.274, 0.422), based on 1,000 bootstrap replications. The proximity of this estimate to those obtained from the current status data underscores the validity and practical utility of our methods, even under severe censoring.
Finally, using the pseudo-value approach, we also tested the mean effect of key prognostic factors, including age, tumor size ($\leq 2$cm, 2-5cm, >5>5cm), nodal status (negative vs. positive), type of surgery (mastectomy with radiation therapy, mastectomy without radiation therapy, or breast conserving), perioperative chemotherapy (yes/no) and adjuvant chemotherapy (yes/no). Initial estimates for the pseudo-value regression were obtained using either the FRE or PLE method. As summarized by Putter et al. [25], these prognostic factors provide important context for modeling progression. Notably, breast-conserving surgery, compared to mastectomy with radiation therapy, was significantly associated with an increased probability of transitioning to the distant metastasis state following a prior visit to the local recurrence state (FRE $p$-value: 0.022; PLE $p$-value: 0.018).
@@ 9 Discussion
This paper carefully employs techniques from the framework of competing risks to construct nonparametric estimators of state entry probabilities and state entry distributions, predicated on the prior occupation of a given state, utilizing current status multistate data. The severe censoring of the event history data posed by a single observation of individuals at different random inspection times brings about additional complexities when attempting to construct nonparametric estimates of temporal functions within a multistate model. Such data are frequently encountered in low-resource settings, where limited follow-up data make it difficult to track disease progression comprehensively. In such scenarios, identifying patients at higher risk of advancing to more severe disease stages is crucial for prioritizing interventions and optimizing resource allocation.
We tackle this problem by harnessing ideas from nonparametric regression modeling to formulate estimators of the counting process and at-risk sets that serve as foundational elements in constructing estimators of the probability functions. To estimate the conditional quantities in the context of current status data, we employed the fraction-risk estimation method that was only available for right-censored data, and we also introduced a novel method based on product-limit estimation. Using data from a real-world application and simulation studies for illustrative purposes, we show that both estimation methods yield valid and comparable results.
In practice, each estimation method has its benefits and challenges. Using the PLE, estimation of $\Psi_{k\mid j}$ depends on estimated marginal state-occupation probabilities for successive states. Consequently, any estimation error in the “starting” marginal SOPs can propagate forward and affect the final estimate. For the FRE approach, if the true fractional weights were known, it would be optimal; in practice, however, these weights must also be estimated recursively, so estimation error may propagate there as well. Computationally, the PLE is easier to implement once the artificial states are specified and marginal SOPs are available from standard software [4]. The FRE approach is more involved because it requires calculating fractional weights and combining information across a sequence of previous states. Nonetheless, in our simulations, FRE generally performed better, showing smaller bias in smaller samples and for transitions that occur deeper in the multistate tree.
The smoothed-bootstrap pointwise confidence intervals exhibit stable and reliable coverage across a wide range of settings, with slightly wider intervals observed for targets involving states deep in the progressive tree and at later times, where current-status sampling provides less information and the effective sample size is smaller. The PLE intervals are typically wider than those from FRE, reflecting additional variability induced by the ratio-of-marginals construction. For a few early time points, mild undercoverage can occur due to finite-sample skewness near the edge of the time range. Overall, the results highlight that current-status inference for late-state probabilities requires adequate inspection-time support and typically larger sample sizes; further improvements may be obtained by time-adaptive smoothing or more targeted bootstrap calibration.
The intuitive approaches we propose provide a more direct route for analyzing the severely censored event data since they utilize the averaging principle rather than maximizing a likelihood function. Suppose we consider a less complicated setting (survival model with two states); the nonparametric maximum likelihood estimator (NPMLE) of the survival function has undesirable properties since it may be undefined over a set of regions. For the general multistate model with current status data, an NPMLE approach would be more problematic.
@@
Appendix A Nonparametric estimators for a multistate system
Recall that $U_{jj^{\prime}}$ denotes the (unobserved) transition time at which an individual transitions from $j$ to $j^{\prime}$. To estimate temporal functions of the multistate system, two essential functionals are the counting process denoted by $N_{jj^{\prime}}(t)$ and the at-risk process denoted by $Y_{j}(t)$. The counting process $N_{jj^{\prime}}(t)$ counts the number of transitions from state $j$ to $j^{\prime}$ in the time interval $[0,t)$. On the other hand, the at-risk process $Y_{j}(t)$ represents the number of individuals who are at risk of transitioning out of state $j$ at time $t$. These processes are expressed as follows
$Y_{j}(t)$ $=\sum^{n}_{i=1}\mathcal{I}(S_{i}(t-)=j),$
where $\mathcal{I}(\cdot)$ denotes the indicator function and S(t)S(t-) is the state occupied just before time $t$. By the law of large numbers,
$n^{-1}N_{jj^{\prime}}(t)\xrightarrow{P}\text{Pr}(U_{jj^{\prime}}\leq t),$
Using the theory of nonparametric regression, Datta and Sundaram [10] proposed nonparametric estimators of $N_{jj^{\prime}}(t)$ and $Y_{j}(t)$ under this context of current status data. For $N_{jj^{\prime}}(t)$, they proposed a two-step estimation approach where first, isotonic regression via the pooled adjacent violators (PAV) algorithm [5] is performed based on the pairs $(C_{i},\ \mathcal{I}(U_{i,jj^{\prime}}\leq C_{i}))$. Let $\widehat{N}^{P}_{jj^{\prime}}(t)$ denote the fitted responses from the isotonic regression. Then secondly, kernel smoothing [21, 23, 32] with a log-concave density, $K>0$ is performed to remove long flat parts of $\widehat{N}^{P}_{jj^{\prime}}(t)$. The resulting estimator of $N_{jj^{\prime}}(t)$ is given by
where $K_{h}(\cdot)=h^{-1}K(\cdot/h)$ with a bandwidth sequence $0<h=h(n)\downarrow 0$ chosen by the criteria specified by Wand and Jones[31] (see Section A.1). Moreover, to estimate $\text{Pr}(X(t-)=j)$, which represents the probability limit of $n^{-1}Y_{j}(t)$, Datta and Sundaram [10] employed the kernel smoothing approach. Accordingly, the estimation of $Y_{j}(t)$ is given by:
where it is assumed that $S_{i}(C_{i}-)=S_{i}(C_{i})$ with probability 1. The estimators in (10) and (11) allow for inference in multistate models with current status data. Datta and Sundaram [10] further demonstrated their utility in estimating key marginal quantities like transition and state occupation probabilities, based on the product-limit method of Andersen et al. [1].
where $\widehat{\pi}_{j^{\prime}}(0)=\{\widehat{Y}_{j^{\prime}}(0+)\}/\{\sum^{Q}_{j=1}\widehat{Y}_{j}(0+)\}$, and $\widehat{P}_{j^{\prime}j}(0,t)$ is the $j^{\prime}j$th element of $\widehat{\boldsymbol{P}}(0,t)$. Datta and Satten [9] demonstrated that (14) provides a consistent estimator for the state occupation probability, even in non-Markov models. This finding allows us to estimate the conditional state occupation probability without the Markov assumption using a product-limit approach similar to the Aalen-Johansen calculation.
A.1 Bandwidth selection
Following earlier work on nonparametric estimation for multistate current-status data [10, 7], we adopt a data-driven approach for selecting the bandwidths used in the kernel-smoothing step of our estimators and in the smoothed bootstrap. For consistency of the kernel estimators, the bandwidth $h$ must satisfy the usual conditions: $h\rightarrow 0$, $nh^{2}\rightarrow\infty$, $nh^{r}\rightarrow 0$, for some $r>2$; which balance bias and variance. Datta et al. [7] recommend using a common bandwidth based on the design distribution of the inspection times. In particular, each smoothed process is obtained using a normal kernel with a bandwidth chosen to be asymptotically $L_{2}$-optimal for estimating the density of the inspection times $C_{i}$, following the direct plug-in described in Wand and Jones [31, Section 3.2]. Practically, this bandwidth can be computed using the dpik function in the KernSmooth R package.
In the simulation studies, this procedure yielded bandwidths in the range 0.34–0.59 on the simulated time scales. In the real-data application, where inspection times ranged from 0.83 to 6755.93 days, the selected bandwidth was $h=246.8$ days ($\approx 0.68$ years).
Appendix B Testing covariate effects
We aim to test the association between covariates and a temporal function, $\theta(t)$, by using pseudo-values as response variables in regression models.
Let $\boldsymbol{Y}_{i}=\{Y_{i}(t_{1}),\ldots,Y_{i}(t_{r})\}$ denote the vector of pseudo-values calculated at time points $(t_{1},\ldots,t_{r})$. To investigate the relationship between covariates and the temporal function, we assume the following model:
$g(\theta(t_{r}\mid\mathbf{Z}_{i}))=\alpha_{r}+\boldsymbol{\zeta}\mathbf{Z}_{i},$ (15)
where $g(\cdot)$ is an appropriate link function, $\mathbf{Z}_{i}$ is the vector of baseline covariates, and $\alpha_{r}$ allows a unique intercept for each time point $t_{r}$. The vector $\boldsymbol{\zeta}$ quantifies the effects of covariates on the temporal function. The pseudo-value regression model (15) is used as a working marginal model for $\theta(t)$. It need not coincide with the full data-generating mechanism; inference focuses on testing whether covariates have any effect on $\theta(t)$ in this marginal sense. Typically, pseudo-values are recommended to be computed at a minimum of five equally spaced time points spanning the range of event times [16, 3, 4].
We utilize the generalized estimating equation [20] to estimate $\boldsymbol{\beta}=(\alpha_{1},\ldots,\alpha_{r},\boldsymbol{\zeta})$. We solve the following estimating equation
$\sum_{i=1}^{m}\boldsymbol{U}_{i}(\boldsymbol{\beta},\alpha)=\boldsymbol{0},$
where
$\boldsymbol{U}_{i}(\boldsymbol{\beta},\alpha)$ $=\boldsymbol{D}_{i}^{T}\boldsymbol{V}_{i}^{-1}(\boldsymbol{Y}_{i}-g^{-1}(\boldsymbol{\beta}\mathbf{Z}_{i})),$
$\boldsymbol{D}_{i}$ $=\frac{\partial g^{-1}(\boldsymbol{\beta}\mathbf{Z}_{i})}{\partial\boldsymbol{\beta}},$
and $\boldsymbol{V}_{i}=\sigma^{2}\boldsymbol{R}_{i}(\alpha)$ is the working covariance matrix with correlation structure $\boldsymbol{R}_{i}(\alpha)$ and dispersion parameter $\sigma^{2}$.
We estimate the variance of $\widehat{\boldsymbol{\beta}}$ using the robust sandwich estimator:
$\widehat{\text{Var}}(\widehat{\boldsymbol{\beta}})=\boldsymbol{B}^{-1}\boldsymbol{M}\boldsymbol{B}^{-1},$
where
$\boldsymbol{B}$ $=\sum_{i=1}^{m}\boldsymbol{D}_{i}^{T}\widehat{\boldsymbol{V}}_{i}^{-1}\boldsymbol{D}_{i},$
$\boldsymbol{M}$ $=\sum_{i=1}^{m}\boldsymbol{U}_{i}(\widehat{\boldsymbol{\beta}},\widehat{\alpha})\boldsymbol{U}_{i}(\widehat{\boldsymbol{\beta}},\widehat{\alpha})^{T},$
[…]